# Boosting Neural Representations for Videos with a Conditional Decoder

### 08 Mar 2024 v2

@@ Abstract
Implicit neural representations (INRs) have emerged as a promising approach for video storage and processing, showing remarkable versatility across various video tasks. However, existing methods often fail to fully leverage their representation capabilities, primarily due to inadequate alignment of intermediate features during target frame decoding. This paper introduces a universal boosting framework for current implicit video representation approaches. Specifically, we utilize a conditional decoder with a temporal-aware affine transform module, which uses the frame index as a prior condition to effectively align intermediate features with target frames. Besides, we introduce a sinusoidal NeRV-like block to generate diverse intermediate features and achieve a more balanced parameter distribution, thereby enhancing the model’s capacity. With a high-frequency information-preserving reconstruction loss, our approach successfully boosts multiple baseline INRs in the reconstruction quality and convergence speed for video regression, and exhibits superior inpainting and interpolation results. Further, we integrate a consistent entropy minimization technique and develop video codecs based on these boosted INRs. Experiments on the UVG dataset confirm that our enhanced codecs significantly outperform baseline INRs and offer competitive rate-distortion performance compared to traditional and learning-based codecs.
@@ 1 Introduction
Figure 1: Video regression with different model sizes under 300 epochs (top) and different epochs under 1.5M model size (bottom). Our boosted methods with only 300 epochs achieve significantly better reconstruction quality than the corresponding baselines with 4800 epochs.
@@ 3.1 Overview
$\boldsymbol{y}_{t}=E(\boldsymbol{x}_{t};\boldsymbol{\phi}),$ (1)
$\boldsymbol{z}_{t}=M({\rm PE}(t);\boldsymbol{\psi}),$
$\boldsymbol{\hat{x}}_{t}=F(\boldsymbol{y}_{t},\boldsymbol{z}_{t};\boldsymbol{\theta}),$
@@ 3.3 Consistent Entropy Minimization
$\mathcal{L}$ $=\mathcal{L}_{d}+\kappa\mathcal{L}_{r}$ (8)
$=\mathcal{L}_{d}+\kappa{\rm ReLU}(R-R_{target}),$
@@ 4.2 Video Regression
Fig. 1 (top) and Table 2 show the regression performance of various methods on the Bunny and UVG datasets at different scales. It is evident that our boosted methods achieve superior reconstruction quality over the corresponding baselines. As detailed in Table 3, the improvements are consistent across all test videos in the UVG dataset. For instance, compared to the NeRV, E-NeRV, and HNeRV baselines on the ReadySetGo video, our boosted versions exhibit considerable improvements of about 2.11dB, 3.77dB, and 2.59dB, respectively. In Fig. 1 (bottom) and Table 4, we further offer a comparison of regression performance between the boosted versions and baselines on the Bunny and Bosphorus videos across different fitting epochs. Notably, our boosted versions at 300 epochs outperform the baselines at 4800 (for Bunny) and 2400 (for Bosphorus) epochs by a large margin.
@@ 4.3 Video Compression
To assess video compression performance, we follow [7] to train a model for each video, rather than encoding all videos together using a single network as in [5, 1]. After model overfitting, we fine-tune these models using our entropy minimization method for 100 epochs with the initial learning rate $5e{-4}$ and cosine learning rate decay. We empirically find that setting $B_{avg}$ to 4 bits optimizes the RD trade-off. For the baselines, both the three-step compression pipeline of NeRV and our CEM technique are applied for comparison. Besides, we compare our boosted models with traditional codecs (H.264 [46], H.265 [41]) and state-of-the-art (SOTA) learning-based codecs (DCVC [18], DCVC-TCM [35]), where H.264 and H.265 are tested using FFmpeg with the veryslow preset and enabling B frames.
@@
4.4 Video Inpainting
As shown in Table 6 and Fig. 6, both the quantitative and qualitative results indicate that our boosted versions significantly improve the inpainting performance of the original baselines. Under the disperse mask case, we observe an average improvement of 3.92dB, 2.43dB, and 4.2dB for NeRV, E-NeRV, and HNeRV, respectively. Even in the more challenging central mask scenario, our enhanced versions still achieve improvements of 2.17dB, 1.37dB, and 1.93dB.
@@ 4.5 Video Interpolation
We evaluate the video interpolation performance of our boosted models on the UVG dataset. In this experiment, we use the odd-numbered frames of each video as the training set and the even-numbered frames as the test set. Following the approach in [19], we adjust the frequency value $b$ to 1.05 for achieving better interpolation results while maintaining robust regression performance on the training set. The quantitative outcomes are detailed in Table 7, and the qualitative performance can be observed in Fig. 6. These results indicate that our boosted models outperform the baselines in terms of overall interpolated quality.
@@ 4.6 Ablation Study
Figure 6: Visualization comparison of disperse inpainting, central inpainting, and interpolation on the DAVIS validation and UVG datasets. The top row displays the ground truth, followed by our boosted results in (a, c, e) and the baseline results in (b, d, f). The red and blue numbers indicate the PSNR values for the respective colored patches.
@@
Feature Modulation. We conduct two sets of ablation studies on Bunny to evaluate the effectiveness of our TAT module. In the first variant (V1), the TAT residual blocks are removed from our boosted INR models. As Table 8 shows, this variant experiences an average drop of 2.97dB, implying the significance of using temporal embeddings to modulate intermediate features for accurate frame generation. In the second variant (V2), the TAT block is replaced with an AdaIN module, resulting in an average decrease in PSNR by 1.76dB. It suggests that integrating normalization into the conditional affine transformation limits the overfitting capabilities of INR models to some degree.
NeRV-like Blocks. Table 8 investigates the effect of different activation layers in SNeRV blocks. By replacing the SINE layer with a common GELU layer (Variant V3), we find that the periodic inductive bias of the SINE layer is beneficial to the reconstruction quality. Subsequently, we explore different convolution configurations. As shown in Table 9 and Fig. 7, our SNeRV block achieves a more uniform parameter distribution compared to previous block designs, contributing to the success of our enhanced video INRs.

### 16 Mar 2024 v3

@@ Abstract
Implicit neural representations (INRs) have emerged as a promising approach for video storage and processing, showing remarkable versatility across various video tasks. However, existing methods often fail to fully leverage their representation capabilities, primarily due to inadequate alignment of intermediate features during target frame decoding. This paper introduces a universal boosting framework for current implicit video representation approaches. Specifically, we utilize a conditional decoder with a temporal-aware affine transform module, which uses the frame index as a prior condition to effectively align intermediate features with target frames. Besides, we introduce a sinusoidal NeRV-like block to generate diverse intermediate features and achieve a more balanced parameter distribution, thereby enhancing the model’s capacity. With a high-frequency information-preserving reconstruction loss, our approach successfully boosts multiple baseline INRs in the reconstruction quality and convergence speed for video regression, and exhibits superior inpainting and interpolation results. Further, we integrate a consistent entropy minimization technique and develop video codecs based on these boosted INRs. Experiments on the UVG dataset confirm that our enhanced codecs significantly outperform baseline INRs and offer competitive rate-distortion performance compared to traditional and learning-based codecs. Code is available at https://github.com/Xinjie-Q/Boosting-NeRV.
@@ 1 Introduction
Figure 1: Video regression with different encoding time under 1.5M model size (left) and video compression with different model sizes(right). Our boosted methods achieve significantly better performance than the corresponding baselines.
@@ 3.1 Overview
$\boldsymbol{y}_{t}=E(\boldsymbol{x}_{t};\boldsymbol{\phi}),\boldsymbol{z}_{t}=M({\rm PE}(t);\boldsymbol{\psi}),\boldsymbol{\hat{x}}_{t}=F(\boldsymbol{y}_{t},\boldsymbol{z}_{t};\boldsymbol{\theta})$ (1)
@@ 3.3 Consistent Entropy Minimization
$\mathcal{L}=\mathcal{L}_{d}+\kappa\mathcal{L}_{r}=\mathcal{L}_{d}+\kappa{\rm ReLU}(R-R_{target})$ (8)
@@ 4.2 Video Regression
Table 2 show the regression performance of various methods on the UVG dataset at different scales. It is evident that our boosted methods achieve superior reconstruction quality over the corresponding baselines. As detailed in Table 3, the improvements are consistent across all test videos in the UVG dataset. For instance, compared to the NeRV, E-NeRV, and HNeRV baselines on the ReadySetGo video, our boosted versions exhibit considerable improvements of about 2.11dB, 3.77dB, and 2.59dB, respectively. In Table 4, we offer a comparison of regression performance between the boosted versions and baselines on the Bosphorus video across different fitting epochs. Notably, our boosted versions at 300 epochs outperform the baselines at 2400 epochs. Furthermore, as shown in Fig. 1 (top), our boosted versions at minimal training time surpass the baselines at their maximum training time by a significant margin on the Bunny video, which indicates the superiority of our method in accelerating convergence speed and improving the representation capabilities.
@@ 4.3 Video Compression
To assess video compression performance, we follow [9] to train a model for each video, rather than encoding all videos together using a single network as in [7, 1]. After model overfitting, we fine-tune these models using our entropy minimization method for 100 epochs with the initial learning rate $5e{-4}$ and cosine learning rate decay. We empirically choose $B_{avg}$ as 4 bits to optimize the RD trade-off. For the baselines, both the three-step compression pipeline of NeRV and our CEM technique are applied for comparison. Besides, we compare our boosted models with traditional codecs (H.264 [50], H.265 [45]) and state-of-the-art (SOTA) learning-based codecs (DCVC [22], DCVC-TCM [39]), where H.264 and H.265 are tested using FFmpeg with the veryslow preset and enabling B frames.
@@
4.4 Video Inpainting
As shown in Table 6, our boosted versions significantly improve the inpainting performance of the original baselines. Under the disperse mask case, we observe an average improvement of 3.92dB, 2.43dB, and 4.2dB for NeRV, E-NeRV, and HNeRV, respectively. Even in the more challenging central mask scenario, our enhanced versions still achieve improvements of 2.17dB, 1.37dB, and 1.93dB.
@@ 4.5 Video Interpolation
We evaluate the video interpolation performance of our boosted models on the UVG dataset. In this experiment, we use the odd-numbered frames of each video as the training set and the even-numbered frames as the test set. Following the approach in [23], we adjust the frequency value $b$ to 1.05 for achieving better interpolation results while maintaining robust regression performance on the training set. The quantitative outcomes are detailed in Table 7, These results indicate that our boosted models outperform the baselines in terms of overall interpolated quality.
@@ 4.6 Ablation Study
(V10) w/ L1+SSIM+freq. 36.99 40.05 41.01
@@
Feature Modulation. We conduct a sets of ablation studies on Bunny to evaluate the effectiveness of our TAT module. In the first variant (V1), the TAT residual blocks are removed from our boosted INR models. As Table 8 shows, this variant experiences an average drop of 2.97dB, implying the significance of using temporal embeddings to modulate intermediate features for accurate frame generation. In the second variant (V2), the TAT block is replaced with an AdaIN [17] module, resulting in an average decrease in PSNR by 1.76dB and 1.19dB. It suggests that integrating normalization into the conditional affine transformation limits the overfitting capabilities of INR models to some degree. Furthermore, we compare the spatially-adaptive fusion (SAF) [15] block. The results in variant V3 imply the superior temporal alignment capability of our TAT module.
NeRV-like Blocks. Table Table 8 investigates the effect of activation layers in SNeRV blocks. By replacing the SINE layer with a common GELU layer (Variant V4), we find that the periodic inductive bias of the SINE layer is beneficial to the reconstruction quality. Subsequently, we explore different convolution configurations. As shown in Table 9 and Fig. 6, integrating SINE in the NeRV, E-NeRV, and FFNeRV blocks results in a significant performance improvement of about 0.7dB. However, only modest improvements are observed in the HNeRV and SNeRV modules. This is attributed to the smaller STD values and more evenly distributed parameters within the HNeRV and SNeRV modules, helping the layers near the output to have sufficient capacity to store high-resolution video content and details. Thus, these modules efficiently fit videos without requiring elaborate feature extraction.